Circular No. 8022

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jf Further to IAUC 8019, W. M. Wood-Vasey et al. report the discovery

of an apparent supernova (mag 19.7) on unfiltered NEAT images taken with the Palomar Schmidt 1.2-m telescope on Oct. 30 UT. Confirming NEAT images were obtained with the Haleakala 1.2-m telescope on Nov. 1 (mag 19.4) and 6 (mag 18.8) and with the Palomar telescope on Nov. 12 (mag 17.8) and 22 (mag 18.7). The new object is located at α = 2^{h}34^{m}35^{s}.43, δ = +12^{o}17′25′′.4 (equinox 2000.0), which is 3′′ east and 1′′ south of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT images taken on Oct. 26 (limiting mag 21.0 at S/N = 3).

SUPERNOVA 2002jg IN NGC 7253 M. Ganeshalingam and W. Li, University of California at Berkeley; and

M. Schwartz, Cottage Grove, OR, report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT images taken on Nov. 23.2 (mag ∼ 17.0) and 24.2 UT (mag ∼ 16.8), and on 0.8-m Tenagra II automatic telescope images taken on Nov. 24.1 (mag ∼ 16.8) and 25.2 (mag ∼ 16.6). SN 2002jg is located at α = 22^{h}19^{m}28^{s}.80, δ = +29^{o}23′04′′.5 (equinox 2000.0), which is 20′′.0 west and 13′′.1 south of the brightest southeast nucleus of the interacting triple-galaxy system NGC 7253. A KAIT image taken on Nov. 18.2 showed nothing at this position (limiting mag ∼ 19.5).

COMETS C/1999 M4, C/2002 U3, C/2002 U4, C/2002 U5 (SOHO) Further to IAUC 8000, D. Hammer reports measurements for Kreutz

sungrazing comets found by M. Oates (C/1999 M4), S. Hönig (C/2002 U3, C/2002 U4), and R. Kracht (C/2002 U5) on SOHO website C2 coronagraph images (with C/1999 M4 and C/2002 U5 also visible in C3 images). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 M4 June 24.071  06^{h}02.^{m}0  +21^{o}15′  2002-V79
Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 U3 Oct. 17.371  13^{h}20.^{m}4  −10^{o}02′  2002-V79
C/2002 U4  22.893  13 42.0  −12 04  2002-W34
C/2002 U5  26.863  13 34.9  −15 31  2002-W34
2002 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green